# Xuezhang Chen  and Shihong Zhang

### v1

@@ 5 Classification theorems
$v(x,t)=\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}f_{0}(y)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tf_{2}(y)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y.$
@@
$|v(x,t)|\lesssim\begin{cases}|X|^{2-\min\{p_{2}c,n\}}&\mathrm{if~{}~{}}p_{2}c\neq n\\ |X|^{2-n}\log|X|&\mathrm{if~{}~{}}p_{2}c=n\end{cases}+\begin{cases}|X|^{-\min\{c,n\}}&\mathrm{if~{}~{}}p_{2}c\neq n\\ |X|^{-n}\log|X|&\mathrm{if~{}~{}}p_{2}c=n\end{cases}\to 0$ (5.8)
@@
By Theorem 1.1, $w:=u-v$ satisfies
$\begin{cases}\displaystyle\Delta^{2}w=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle w(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\Delta w(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
Moreover, by (5.8) we have
@@
where $\delta=\min\left\{\min\{p_{2}c,n\}-2,\min\{c,n\}\right\}/2>0$. By Lemma 2.4 we have
@@
where $c_{3}$ is a nonnegative constant.
@@
$\lim_{t\to 0}\partial_{t}\int_{\mathbb{R}^{n}}\frac{t^{3}}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}u(y,0)\mathrm{d}y=0,$
and the first boundary condition we obtain
@@
Since $u(x,0)=o(1)$ as $|x|\to\infty$, by Lemma 2.1 we have
@@
(1) If $\frac{p_{2}}{p_{1}}=\frac{n+1}{n-1}$, then we apply the classification theorem (see [9] and [24]) to know
@@
By (5.2) we have
$u(x,t)=M_{2}(x,t)+c_{3}t^{3}$
and
@@
$w(x,t)=u(x,t)-\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}},$
which satisfies
$\begin{cases}\displaystyle\Delta^{2}w=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle w(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\Delta w(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (5.11)
with
@@
$u(x,t)=\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}}+c_{3}t^{3}.$
(2) If $\frac{p_{2}}{p_{1}}<\frac{n-1}{n-3}$, then the uniqueness theorem in Y. Y. Li [24] yields $u(x,0)=0$ in $\mathbb{R}^{n}$, thus $u(x,t)$ satisfies the homogeneous boundary value problem (5.11). This together with Lemma 2.4 yields $u(x,t)=c_{3}t^{3}$ in $\mathbb{R}_{+}^{n+1}$ for every nonnegative constant $c_{3}$. $\Box$
@@
Proof of Theorem 1.8. We now consider
$\begin{cases}\displaystyle\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}_{2}^{3}(u)=\frac{(n-1)(n-3)}{2}u^{p_{2}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\frac{(n^{2}-1)(n-3)}{4}u^{p_{3}}&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (5.12)
@@
We first claim that
$P_{2}(x)\equiv c\geq 0\qquad\mathrm{and}\qquad c_{3}=0.$
It is known from (5.14) that $\overline{\Delta}P_{2}(x)$ is a nonnegative constant. With $\tilde{c}_{3}:=c_{3}+\frac{1}{2}\overline{\Delta}P_{2}(x)$ we define
@@
Clearly, $\tilde{u}$ is biharmonic in $\mathbb{R}_{+}^{n+1}$ and $\tilde{u}(x,0)=u(x,0)$. Recall that $\mathscr{B}^{3}_{2}=\partial_{t}^{2}-\overline{\Delta}$ and $\mathscr{B}^{3}_{3}=\partial_{t}\Delta+2\overline{\Delta}\partial_{t}$, then
@@
and $\mathscr{B}^{3}_{2}(\tilde{u})=\mathscr{B}^{3}_{2}(u)$ on $\partial\mathbb{R}_{+}^{n+1}$. This means that $\tilde{u}$ also satisfies the biharmonic boundary value problem (5.12), though it might change sign.
On the other hand, notice that $\tilde{u}(x,t)\geq-C|X|^{3}$ as $|X|\to\infty$. By Lemma 2.1 as $|x|\to\infty$ we have
$-\partial_{t}\tilde{u}(x,0)=\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{\tilde{u}^{p_{2}}(y,0)}{|x-y|^{n-1}}\mathrm{d}y=$$\begin{cases}O(|x|^{1-\min\{p_{2}c,n\}})&\qquad\mathrm{if~{}~{}}p_{2}c\neq n\\ O(|x|^{1-n}\log|x|)&\qquad\mathrm{if~{}~{}}p_{2}c=n\end{cases}$
$=$ $O(|x|^{-d}),$
where $d=\min\{cp_{2}-1,(n-1)/2\}>1$. For $c>\max\{2/p_{2},3/p_{3}\}$, we apply Remark 2.1 to obtain
$\tilde{u}(x,t)=\frac{2}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}\partial_{t}\tilde{u}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y+\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{3}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\mathrm{d}y+t^{2}P_{1}(x).$ (5.16)
By (5.3) and (5.16), we can apply Lemmas 2.1 and 2.2 to conclude that
@@
This implies $\tilde{c}_{3}=0$ and $P_{1}(x)\equiv 0$. Then, (5.16) becomes
@@
$0=\tilde{c}_{3}=c_{3}+\frac{\overline{\Delta}P_{2}(x)}{2},$
so, $c_{3}=0$ and $\overline{\Delta}P_{2}(x)=0$. This together with (5.14) implies $P_{2}(x)\equiv c_{1}\geq 0$. Thus, we prove the above claim.
Going back to (5.3) and (5.3) we now have
@@
Furthermore, letting $t\to 0$ in (5.17) we arrive at
@@
(1) If $p_{3}=\frac{n+3}{n-3}$ and $p_{2}>\frac{2}{n-3}$, then the classification theorem in [9] and [24] gives
@@
In particular, when $p_{2}=p_{2}^{\ast}$, as before, we notice that $(\tilde{u}-U_{\varepsilon,x_{0}})(X)\geq-C|X|$ as $|X|\to\infty$, and apply Remark 2.2 to conclude that $M_{3}=U_{\varepsilon,x_{0}}$ and then
$u(x,t)=U_{\varepsilon,x_{0}}(x,t)+c_{1}t.$
(2) If $p_{3}<\frac{n+3}{n-3}$, then the uniqueness theorem in [24] implies $u(x,0)=0$. This together with (5.3) implies $u(x,t)=c_{1}t$ in $\mathbb{R}_{+}^{n+1}$ for every nonnegative constant $c_{1}$. $\Box$
@@ 6 Applications to biharmonic boundary value problems on balls
Although we focus on the conformal geometry nature for biharmonic boundary value problems, we would like to mention that from a physical point of view, the boundary conditions $(\mathscr{B}_{0}^{3},\mathscr{B}_{1}^{3})$ and $(\mathscr{B}_{0}^{3},\mathscr{B}_{2}^{3})$ have a close connection with the clamped plate and hinged plate model (cf. [15]), respectively. However, the study of other boundary conditions is far less developed.
6.1 One boundary point singularity
Proof of Theorem 1.9. It suffices to give a proof for $i=1$ and $j=3$, since the others are the same in the spirit due to Theorems 1.7 and 1.8.
@@ 6.1 One boundary point singularity
$u(X)=\left(\frac{2}{|X+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}U\circ F.$
@@
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(u)=\mathbb{T}^{3}_{1}u^{\frac{n-1}{n-3}}&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{j}(u)=\mathbb{T}^{3}_{3}u^{\frac{n+3}{n-3}}&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}_{+}^{n+1}.\end{cases}$
By definition of the spherical inversion $F$, we know that $|X+\mathbf{e}_{n+1}||\xi+\mathbf{e}_{n+1}|=2$ and $F(\partial\mathbb{R}_{+}^{n+1})=\partial\mathbb{B}^{n+1}$. By the assumption (1.10) for $U$, we obtain
@@ 6.2 Sharp geometric inequalities
On a compact manifold with boundary, the maximizing problem of the isometric ratio over a scalar-flat conformal class was studied in [22, 21]. With comparison principles in Proposition 4.1, we are motivated to study of the maximizers of the isometric ratio and another sharp geometric inequality over some $Q$-curvature flat conformal class on $(\overline{\mathbb{B}^{n+1}},|\mathrm{d}\xi|^{2})$.
@@
$\mathcal{M}_{1}=$$\{g\in[|\mathrm{d}\xi|^{2}]~{}|~{}Q_{g}\leq 0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}(T_{2}^{3})_{g}\geq 0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\};$
$\mathcal{M}_{2}=$$\{g\in[|\mathrm{d}\xi|^{2}]~{}|~{}Q_{g}\leq 0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}h_{g}\geq 0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\}$
@@
$\mathring{\mathcal{M}}:=$$\{g\in[|\mathrm{d}\xi|^{2}]~{}|~{}Q_{g}=0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}(T_{2}^{3})_{g}=0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\}$
$=$ $\{g\in[|\mathrm{d}\xi|^{2}]~{}|~{}Q_{g}=0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}h_{g}=0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\}$
@@
Let $n\geq 4$. Then for all $g\in\mathcal{M}_{1}\cup\mathcal{M}_{2}$, there hold
@@
Gluck [17, Theorem 1.1] states that the optimal constant $d_{n}$ for the extremal problem (6.4) is attained by
@@
equivalently, modulo of a positive constant, the extremal metric on $\mathbb{S}^{n}$ is a conformal metric $u_{a}^{4/(n-3)}g_{\mathbb{S}^{n}}$ induced by some conformal transformation on $\mathbb{S}^{n}$, which can be expressed as
@@
such that $U_{a}^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}\in\mathring{\mathcal{M}}$ is an extremal metric with $U_{a}=u_{a}$ on $\mathbb{S}^{n}$. Indeed, the trick in authors [11, Section 6] can be applied to derive an explicit representation formula of
@@
Dou-Zhu [12, Theorem 1.4] states that every maximizer for (6.7) has the form of
@@
Combining the above with authors [11, Section 6], we conclude from (6.6) that, modulo a positive constant,

### v2

@@ 5 Classification theorems
$v(x,t)=\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}f_{0}(y)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tf_{2}(y)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y$
@@
$|v(x,t)|\lesssim\begin{cases}|X|^{-\min\{c,n\}}&\mathrm{if~{}~{}}c\neq n\\ |X|^{-n}\log|X|&\mathrm{if~{}~{}}c=n\end{cases}+\begin{cases}|X|^{2-\min\{p_{2}c,n\}}&\mathrm{if~{}~{}}p_{2}c\neq n\\ |X|^{2-n}\log|X|&\mathrm{if~{}~{}}p_{2}c=n\end{cases}\to 0$ (5.8)
@@
By the assumption and Theorem 1.1, $w:=u-v$ satisfies
$\begin{cases}\displaystyle\Delta^{2}w=0\qquad\qquad\qquad&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\quad~{}w=0&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle~{}~{}\Delta w=0&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
Moreover, using the nonnegativity of $u$ and (5.8) we have
@@
where $\delta=\min\left\{\min\{p_{2}c,n\}-2,\min\{c,n\}\right\}/2>0$. Then we apply Lemma 2.4 to obtain
@@
where $c_{3}$ is a nonnegative constant and $\liminf_{|x|\to\infty}P_{2}(x)\geq 0.$
@@
$\lim_{t\to 0}\partial_{t}\int_{\mathbb{R}^{n}}\frac{t^{3}u(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y=0,$
together with the first boundary condition we obtain
@@
Using the assumption that $u(x,0)=o(1)$ as $|x|\to\infty$ and Lemma 2.1 we have
@@
(1) If $\frac{p_{2}}{p_{1}}=\frac{n+1}{n-1}$, then the classification theorem (e.g., [12] and [30]) yields
@@
Going back to (5.2) we conclude that
$u(x,t)=M_{2}(x,t)+c_{3}t^{3},$
where
@@
$w(x,t)=u(x,t)-U_{x_{0},\varepsilon}(x,t).$
Then $w$ solves
$\begin{cases}\displaystyle\Delta^{2}w=0\qquad\qquad\qquad&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\quad~{}w=0&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle~{}~{}\Delta w=0&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (5.11)
with some $C\in\mathbb{R}_{+}$ such that
@@
$u(x,t)=U_{x_{0},\varepsilon}(x,t)+c_{3}t^{3}.$
(2) If $\frac{p_{2}}{p_{1}}<\frac{n-1}{n-3}$, then the uniqueness theorem in Y. Y. Li [30] yields $u(x,0)=0$ in $\mathbb{R}^{n}$, so $u$ satisfies problem (5.11). This together with Lemma 2.4 yields $u(x,t)=c_{3}t^{3}$ in $\mathbb{R}_{+}^{n+1}$ for every nonnegative constant $c_{3}$. $\Box$
@@
Proof of Theorem 1.6. Recall that $u$ solves
$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\quad~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}_{2}^{3}(u)=\mathbb{T}^{3}_{2}u^{p_{2}}&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\mathbb{T}^{3}_{3}u^{p_{3}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (5.12)
@@
We claim that there is some nonnegative constant $c$ such that
$P_{2}(x)\equiv c\qquad\mathrm{~{}~{}and~{}~{}}\qquad c_{3}=0.$
To this end, (5.14) yields that $\overline{\Delta}P_{2}(x)$ must be a nonnegative constant. With $\tilde{c}_{3}:=c_{3}+\frac{1}{2}\overline{\Delta}P_{2}(x)$ we introduce
@@
Clearly, $\tilde{u}$ is biharmonic in $\mathbb{R}_{+}^{n+1}$ and $\tilde{u}(x,0)=u(x,0)$. Recall that $\mathscr{B}^{3}_{2}=\partial_{t}^{2}-\overline{\Delta}$ and $\mathscr{B}^{3}_{3}=\partial_{t}\Delta+2\overline{\Delta}\partial_{t}$, then $\tilde{u}$ and $u$ satisfy the same boundary conditions, more explicitly,
@@
and $\mathscr{B}^{3}_{2}(\tilde{u})=\mathscr{B}^{3}_{2}(u)$ on $\partial\mathbb{R}_{+}^{n+1}$. This means that $\tilde{u}$ also satisfies problem (5.12), though it might change sign in $\mathbb{R}_{+}^{n+1}$.
On the other hand, notice that there is some $C\in\mathbb{R}_{+}$ such that $\tilde{u}(x,t)\geq-C|X|^{3}$ as $|X|\to\infty$. By Lemma 2.1 we have
$-\partial_{t}\tilde{u}(x,0)=$ $\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{\tilde{u}^{p_{2}}(y,0)}{|x-y|^{n-1}}\mathrm{d}y$
$=$ $\begin{cases}O(|x|^{1-\min\{p_{2}c,n\}})&\qquad\mathrm{if~{}~{}}p_{2}c\neq n\\ O(|x|^{1-n}\log|x|)&\qquad\mathrm{if~{}~{}}p_{2}c=n\end{cases}$
$=$ $O(|x|^{-d})\qquad\qquad\mathrm{~{}~{}as~{}~{}}|x|\to\infty,$
where $d=\min\{cp_{2}-1,(n-1)/2\}>1$. As $c>\max\{2/p_{2},3/p_{3}\}$ by assumption, we apply Remark 2.1 to obtain
$\tilde{u}(x,t)$
$=$ $\frac{2}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}\partial_{t}\tilde{u}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y+\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{3}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\mathrm{d}y+t^{2}P_{1}(x).$ (5.16)
Combining (5.3) and (5.3), we can apply Lemmas 2.1 and 2.2 to conclude that
@@
This implies $\tilde{c}_{3}=0$ and $P_{1}(x)\equiv 0$. Thus, (5.3) becomes
@@
$0=\tilde{c}_{3}=c_{3}+\frac{1}{2}\overline{\Delta}P_{2}(x),$
so, $c_{3}=0$ and $\overline{\Delta}P_{2}(x)=0$. This together with (5.14) implies $P_{2}(x)\equiv c_{1}\geq 0$. Hence, we prove the above claim.
Going back to (5.3) and (5.3) we have
@@
Furthermore, letting $t\to 0$ in (5.17), by Lemma 2.1 we arrive at
@@
(1) If $p_{3}=\frac{n+3}{n-3}$ and $p_{2}>\frac{2}{n-3}$, then the classification theorem (e.g., [30] and [12]) gives
@@
In particular, when $p_{2}=p_{2}^{\ast}$, notice that there is some $C\in\mathbb{R}_{+}$ such that $(\tilde{u}-U_{x_{0},\varepsilon})(X)\geq-C|X|$ as $|X|\to\infty$. Then, we apply Remark 2.2 to conclude that $M_{3}=U_{x_{0},\varepsilon}$ and then
$u(x,t)=U_{x_{0},\varepsilon}(x,t)+c_{1}t.$
(2) If $p_{3}<\frac{n+3}{n-3}$, then the uniqueness theorem (e.g., [30]) yields $u(x,0)=0$. This together with (5.3) follows that $u(x,t)=c_{1}t$ in $\mathbb{R}_{+}^{n+1}$ for every nonnegative constant $c_{1}$. $\Box$
@@ 6 Applications to biharmonic boundary value problems on balls
Although we focus on the conformal geometry nature for biharmonic BVPs, we would like to mention that from a physical viewpoint, the boundary conditions $(\mathscr{B}_{0}^{3},\mathscr{B}_{1}^{3})$ and $(\mathscr{B}_{0}^{3},\mathscr{B}_{2}^{3})$ have a close connection with the clamped plate and hinged plate model (cf. [20]), respectively. However, the study of biharmonic equation coupled with other boundary conditions is far less developed.
6.1 Solutions with single boundary point singularity
Proof of Theorem 1.7. It suffices to give the proof for $i=1$ and $j=3$, since the others are the same in the spirit thanks to Theorems 1.5 and 1.6.
@@ 6.1 Solutions with single boundary point singularity
$u(X)=\left(\frac{2}{|X+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}(U\circ F)(X).$
@@
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{1}(u)=\mathbb{T}^{3}_{1}u^{\frac{n-1}{n-3}}&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{3}(u)=\mathbb{T}^{3}_{3}u^{\frac{n+3}{n-3}}&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}_{+}^{n+1}.\end{cases}$
By definition of the conformal map $F$, we know that $|X+\mathbf{e}_{n+1}||\xi+\mathbf{e}_{n+1}|=2$ and $F(\partial\mathbb{R}_{+}^{n+1})=\partial\mathbb{B}^{n+1}$. By the assumption (1.11) for $U$, we obtain
@@ 6.2 Sharp geometric inequalities
On a smooth compact manifold with boundary, the maximizing problem of the isometric ratio over a scalar-flat conformal class was studied in [27, 26]. With comparison principles in Proposition 4.1, we are motivated to study of the maximizers of the isometric ratio and another sharp geometric inequality over some $Q$-curvature flat conformal class of the flat metric $[|\mathrm{d}\xi|^{2}]$ on $\overline{\mathbb{B}^{n+1}}$.
@@
$\mathcal{M}_{1}=$$\{g\in[|\mathrm{d}\xi|^{2}];Q_{g}\leq 0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}(T_{2}^{3})_{g}\geq 0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\};$
$\mathcal{M}_{2}=$$\{g\in[|\mathrm{d}\xi|^{2}];Q_{g}\leq 0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}h_{g}\geq 0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\}$
@@
$\mathring{\mathcal{M}}:=$$\{g\in[|\mathrm{d}\xi|^{2}];Q_{g}=0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}(T_{2}^{3})_{g}=0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\}$
$=$ $\{g\in[|\mathrm{d}\xi|^{2}];Q_{g}=0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}h_{g}=0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\}$
@@
Let $n\geq 4$. Then for all $g\in\mathcal{M}_{1}\cup\mathcal{M}_{2}$,
@@
Gluck [22, Theorem 1.1] states that the optimal constant $d_{n}$ for the extremal problem (6.4) is attained by
@@
equivalently, modulo of a positive constant, the extremal metric is a conformal metric $u_{a}^{4/(n-3)}g_{\mathbb{S}^{n}}$ induced by some conformal transformation on $\mathbb{S}^{n}$, which can be expressed as
@@
such that $U_{a}^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}\in\mathring{\mathcal{M}}$ is an extremal metric with $U_{a}=u_{a}$ on $\mathbb{S}^{n}$. Indeed, the trick in authors [15, Section 6] can be applied to derive an explicit representation formula of
@@
Dou-Zhu [16, Theorem 1.4] states that every maximizer for (6.7) has the form of
@@
Combining the above discussion and the authors [15, Section 6], we conclude from (6.6) that modulo a positive constant,
@@
7 Perspectives
Similar to the boundary Yamabe problem (see for example Chen-Sun [14] and Escobar [17, 18]), it is natural to ask whether the round metric on a geodesic ball $\Sigma_{r}^{n+1}$ with radius $r\in(0,\pi)$ in $\mathbb{S}^{n+1}$ is the unique conformal metric (up to a conformal diffeomorphism) of constant $Q$-curvature and constant boundary $T$-curvature under certain geometric constraints on the pair of $T$-curvature constants.
Consider $\Sigma_{r}^{n+1}$ equipped with the round metric $g_{c}:=g_{\mathbb{S}^{n+1}}$. Using the precise formulae of $T$-curvatures in [9, p.296] and $Q$-curvature in [15], we directly compute that $Q_{g_{\mathrm{c}}}=Q_{\mathbb{S}^{n+1}}=\frac{(n+1)((n+1)^{2}-4)}{8}$ and
$(T^{3}_{1})_{g_{\mathrm{c}}}=$ $h_{g_{\mathrm{c}}}=\cot r,$
$(T^{3}_{2})_{g_{\mathrm{c}}}=$ $(n-1)\big{(}\frac{1}{2}+h_{g_{\mathrm{c}}}^{2}\big{)},\qquad(T^{3}_{3})_{g_{\mathrm{c}}}=\frac{n^{2}-1}{2}h_{g_{\mathrm{c}}}\big{(}\frac{3}{2}+h_{g_{\mathrm{c}}}^{2}\big{)}.$ (7.1)
Notice that $\mathbb{B}^{n+1}$ is conformal equivalent to $\Sigma_{r}^{n+1}$, see for example, the first author and Sun [14, pp.8-9]. Based on (7), we can transfer the above problem into the following conjecture.
Conjecture..
For $n\geq 4$ and $i\in\{1,2\}$, let $U\in C^{4}(\overline{\mathbb{B}^{n+1}})$ be a positive smooth solution to
$\begin{cases}\displaystyle\quad\Delta^{2}U=\frac{n-3}{2}Q_{\mathbb{S}^{n+1}}U^{p^{*}}&\mathrm{~{}~{}in~{}~{}}\quad~{}~{}\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(U)=c_{i}\mathbb{T}^{3}_{i}U^{p^{*}_{i}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad~{}~{}\mathbb{S}^{n},\\ \displaystyle\mathscr{B}_{3}^{3}(U)=c_{3}\mathbb{T}^{3}_{3}U^{p^{*}_{3}}&\mathrm{~{}~{}on~{}~{}}\quad~{}~{}\mathbb{S}^{n},\end{cases}$ (7.2)
with $c_{k}\in\mathbb{R},k\in\{1,2,3\}$ and
$\begin{cases}\displaystyle c_{3}=\tfrac{3}{2}c_{1}+c_{1}^{3}\qquad&\mathrm{for}\qquad i=1,\\ \displaystyle c_{2}\geq\tfrac{1}{2},\quad c_{3}^{2}=(c_{2}-\tfrac{1}{2})(1+c_{2})\qquad&\mathrm{for}\qquad i=2.\end{cases}$
Then $U$ is of the form
$U(\xi)=\left(\frac{2\varepsilon(1-|\xi_{0}|^{2})}{(\varepsilon^{2}+1)(|\xi_{0}|^{2}|\xi|^{2}-2\xi_{0}\cdot\xi+1)+(1-|\xi_{0}|^{2})(|\xi|^{2}-1)}\right)^{\frac{n-3}{2}}$ (7.3)
for some $\xi_{0}\in\mathbb{B}^{n+1}$ and
$c_{1}=\cot r,\quad c_{1}=\frac{\varepsilon^{2}-1}{2\varepsilon}\qquad\mathrm{~{}~{}for~{}~{}}\qquad i=1,$
$c_{2}=\frac{\varepsilon^{4}+1}{4\varepsilon^{2}}\qquad\mathrm{~{}~{}for~{}~{}}\qquad i=2.$
This conjecture can be regarded as an extension of Corollary 1.1 to the constant $Q$-curvature and constant $T$-curvature equations on $\mathbb{B}^{n+1}$. Moreover, we can prove a uniqueness theorem of positive solutions to a general 4-th order ODE including the BVP (7.2) for radial solutions. See Theorem B.1, whose proof is left to Appendix B. On the other hand, it is direct to verify that the radial function $\left(\frac{2\varepsilon}{\varepsilon^{2}+|\xi|^{2}}\right)^{\frac{n-3}{2}}$ corresponding to $\xi_{0}=0$ in (7.3) satisfies (7.2). Hence we confirm the above conjecture for all radial solutions.
We hope that the biharmonic Poisson kernel and Green function will come into play in the resolution of the above conjecture.